# XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge

### v2

@@ Abstract
Deep learning for human sensing on edge systems presents significant potential for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and poor adaptability in practice. In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design. XTransfer flexibly uses single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
@@ 1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a; Ystgaard et al., 2023). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed & Farid, 2018). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020). However, existing DL solutions are often data- and resource-intensive for training and deployment (Dhar et al., 2021), posing significant challenges for data collection and edge deployment in human sensing (Sobin, 2020; Liu et al., 2019).
@@
Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane & Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022). Manual annotation of large volumes of sensor data is labor-intensive and incurs high costs (Song et al., 2023). Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
@@
Prior research has explored Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane & Ploetz, 2024), and their combination (i.e., cross-domain FSL) (Thukral et al., 2025; Yin et al., 2024) to notably reduce data collection costs by adapting pre-trained models to sensing applications. However, they still rely on large-scale labeled datasets from the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift and poor adaptability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings. Recent advances in multi-modal learning (Yang et al., 2023; Chen & Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), demonstrate the feasibility of knowledge transfer across modalities, but they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge. These limitations motivate the need for a scalable and adaptable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
@@
In practice, however, addressing this need is uniquely challenging. Due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022). Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019). While methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle & Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
Figure 1: Overview. XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control). LWS control first segments source models into layers and operates layer-wise search across pools. At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest. These layers are incrementally recombined during the search, resulting in a compact model for enabling human sensing at the edge. Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
@@
Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer. It flexibly leverages public pre-trained models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few sensor data, significantly reducing the costs of large-scale data collection and training from scratch. Its resource-efficient design further optimizes edge deployment, enabling both scalability and adaptability to diverse human sensing applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained models under cross-modality FSL settings (see Appendix A), resulting in compact models. To mitigate modality shift and prevent overfitting, it enables model repairing using a novel Splice-Repair-Removal (SRR) pipeline (see Section 4). SRR pipeline transforms a few sensor data to broadly adapt pre-trained layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (see Section 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, LWS control discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (White et al., 2023). To further accelerate the search in multi-source settings, a pre-search check strategy is designed to efficiently avoid unnecessary repairing.
The main contributions are summarized as: 1) a first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems, as shown in Figure 1; 2) a novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities; 3) a novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner to create compact models; and 4) extensive experimental evaluations (see Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on 8 public datasets with corresponding pre-trained source models across diverse modalities (e.g., image, text, audio, and sensing) and 7 target datasets (Table 2) using 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (see Section 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while reducing layer search time by 2.1 to 4 times and substantially lowering the costs of data collection, model training, and edge deployment, demonstrating broad adaptability.
@@
Figure 2: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting (see Appendix A).. (b) shows the average similarity and FSL difficulty across all sensing datasets to source modalities (e.g., Image, Text, Sensing) using default reshaping (see Appendix B.1). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Section 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability. (c) shows model pruning bottlenecks.
@@ 2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based objectives (Zhao et al., 2021; Yue et al., 2021b), whereas multi-source methods leverage the Wisdom of the Crowd principle through unsupervised source distillation (Zhao et al., 2020; Yue et al., 2021a) or maximal correlation analysis (Lee et al., 2019). Recent efforts have also explored fine-tuning LLMs (e.g., GPT2 (Zhou et al., 2023)) for time-series analysis. However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings. LLM-based data augmentation (Leng et al., 2024; 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus. Existing attempts on human sensing tasks (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training with limited adaptability. In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Frankle & Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model structure optimization (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures while strengthening layer-wise dependence, by efficiently searching and recombining layers of interest from pre-trained source models at scale. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
@@ 3 Preliminary
Figure 3: Design insights. (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift. (b) A notable MMC shift emerges and grows with increasing layer index, i.e., accuracy increases while MMC shift drops in Area A and begins to drop as MMC shift largely grows in Area B. (c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
Insights in cross-modality FSL. We apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights. It uses Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets including 5 human sensing and 2 well-known image datasets (Table 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Figure 2(b).
@@
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (Islam et al., 2021), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (Luo et al., 2022) of each layer as the default metric to quantify layer-wise performance, as it offers a lightweight measure for channel importance. We define a single layer or a dependent layer block (i.e., L-Blocks in Table 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Figure 3(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Figure 3(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (Shahapure & Nicholas, 2020), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts under cross-modality FSL settings, but several technical challenges arise. First, it is hindered by structural and semantic gaps between sources and target data, where classes and semantics often differ. Second, the pre-trained feature space at each layer is highly complex due to the multidimensional and noisy nature of MMC across source models from different modalities. This complexity increases with layer depth, as the strong layer-wise dependence formed during source model training becomes increasingly rigid, making it difficult to repair layer-wise misalignment using only few sensor data. Third, Figure 3(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence. It indicates that not all repaired layers contribute positively, motivating our LWS control design to selectively recombine layers of interest while discarding others. The selection must also account for on-device resource constraints to control layer growth, while the rapidly expanding search space with increasing layers and source models further challenges the layer search.

### v3

@@ Abstract
Deep learning for human sensing on edge systems presents significant potential for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and limited transferability. In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design. XTransfer flexibly uses pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to restructure models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. The results show that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
Human Sensing, Model Transfer, Edge Deployment
@@ 1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed and Farid, 2018). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020). However, existing DL solutions are often data- and resource-intensive for training and deployment (Dhar et al., 2021), posing significant challenges for data collection and edge deployment in human sensing (Sobin, 2020; Liu et al., 2019).
@@
Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane and Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022). Manual annotation of large volumes of sensor data is labor-intensive and incurs high costs (Song et al., 2023). Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
@@
Prior works on Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane and Ploetz, 2024), and cross-domain FSL (Thukral et al., 2025; Yin et al., 2024) reduce data collection costs by adapting pre-trained models to sensing applications. However, they typically rely on large-scale labeled datasets within the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift with limited transferability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings. Recent advances in multi-modal learning (Yang et al., 2023; Chen and Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), further demonstrate the feasibility of knowledge transfer across modalities. However, they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale unlabeled and paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge. These gaps motivate a scalable and adaptable model transfer that can broadly reuse pre-trained models across modalities (e.g., vision or text) for new sensing modalities (e.g., radar or bio-signals) using only few data.
@@
In practice, however, addressing this need is uniquely challenging. Due to substantial differences in data characteristics, such as shape, distribution, and feature representation, between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022). Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a). These highlight the underlying challenge of latent feature distribution misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019). While model structuring methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle and Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
@@
Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer. It flexibly leverages public pre-trained models and transfers knowledge across modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few labeled sensor data with no additional unlabeled data (i.e., cross-modality FSL settings, Appendix A.1), significantly reducing the costs of large-scale data collection and training from scratch. Its resource-efficient design also optimizes both cloud performance and edge deployment, enabling scalable and adaptable method across human sensing applications.
Instead of relying on shared semantic spaces or paired cross-modal data, XTransfer mitigates modality shift by model repairing at the level of layer-wise latent feature distribution misalignment using activation-based statistics. This layer-wise formulation is motivated by our observation (Section 3) that modality shift “damages” intermediate representations unevenly across layers, making global adaptation with few data unstable and prone to overfitting. It is enabled by a Splice–Repair–Removal (SRR) pipeline (Section 4) that (i) splices heterogeneous layer inputs, (ii) repairs layer misalignment via an anchor-based generative transfer module in a reduced-orthogonal feature space, and (iii) removes unnecessary channels after repair. Building on the insight that not all repaired layers contribute, XTransfer further features layer recombining using a Layer-Wise Search (LWS) control (Section 5) that selects and recombines layers of interest (i.e., useful repaired layers) through an NAS-inspired, resource-constrained layer-wise search, incrementally enhancing the repairing and optimizing model structure. To accelerate and stabilize LWS in multi-source settings with large candidate pools, we incorporate a pre-search check strategy coupled with a dynamic search range mechanism to avoid unnecessary repairing, balancing accuracy and efficiency. The main contributions are summarized as:
  • A first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems.
  • A SRR pipeline that mitigates modality shift and prevents overfitting via layer-wise repairing of latent feature distribution misalignment.
  • A LWS control that enables efficient, stable layer-wise search, selectively recombining layers of interest while discarding others for model restructuring.
  • Extensive experimental evaluations (Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on various source and target datasets (Table 2), showing consistent accuracy and efficiency gains.
@@
Figure 1: Preliminary study under cross-modality FSL settings  [^1] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting with LOOCV enabled (details in Appendix A.2).. (a) reveals baseline performance gap. (b) shows average similarity and FSL difficulty across all target sensing datasets in Table 2 to source modalities (e.g., Image, Text, Sensing) using default reshaping (Appendix B.1) and benchmarking (Oh et al., 2022) (Appendix A.2 for details). Two distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes  [^2] The number of classes in the source model should be greater than that in the target. may enhance cross-modality similarity; 3) Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for model transfer. (c) shows significant model accuracy loss using pruning.
@@ 2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based objectives (Zhao et al., 2021; Yue et al., 2021b), whereas multi-source methods leverage the Wisdom of the Crowd principle through unsupervised source distillation (Zhao et al., 2020; Yue et al., 2021a) or maximal correlation analysis (Lee et al., 2019). Recent efforts have also explored fine-tuning LLMs (e.g., GPT2 (Zhou et al., 2023)) for time-series analysis. However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings. LLM-based data augmentation (Leng et al., 2024, 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus. Recent advances introduce foundation models (FM) (Abbaspourazad et al., 2024; Weng et al., 2024) or prompt-based FM adaptation (Li et al., 2025) for human sensing, but these methods assume either a pre-existing target-modality FM or access to large-scale unlabeled and paired data for feature alignment. Existing attempts in human sensing FSL (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training. In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction.
Model structuring for edge deployment. Model structuring techniques such as pruning (Frankle and Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model restructuring (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures while strengthening repaired layer-wise dependence at scale under cross-modality FSL settings. It not only restructures models for streamlined edge deployment but also enhances model repairing.
Figure 2: Design insights. (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift. (b) In area A, accuracy rises as MMC shift stays low, indicating a small latent feature gap. In area B, MMC shift notably increases with layer index, where excessive latent feature deviation begins to reduce accuracy. (c) After repairing, S-score improves but stagnation occurs at certain layers.
@@ 3 Preliminary
Layer-wise analysis. It is widely used to diagnose representation quality and guide model restructuring (Sahoo et al., 2025). We examine how accuracy and latent feature discriminability change across layers. Since effective training relies on discriminative latent representations (Islam et al., 2021), we quantify layer-wise latent features using Mean Magnitude of Channels (MMC) (Luo et al., 2022) as a lightweight activation-based statistics. We also measure class clustering via the Silhouette score (S-score) (Shahapure and Nicholas, 2020) computed on MMC-based representations, which captures inter-/intra-class distances and reflects layer-wise discriminability. We segment the backbone into L-units, defined as a single layer or an inseparable dependent layer block (L-Blocks in Table 3) that cannot be structurally decoupled (e.g., residual blocks in ResNet). The experimental setup is detailed in Appendix A.3, and additional evidence on S-score/accuracy correlation is provided in Appendix A.4. Figure 2(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., impaired layer-wise dependence), leading to misalignment and accuracy drops under few-shot fine-tuning. These degradations correlate with MMC shifts, where larger shifts correspond to larger accuracy drops (Figure 2(b)). We adopt MMC as a modality-agnostic metric as it is defined purely on layer activations without assuming shared semantic spaces or paired data, remaining lightweight and stable under FSL settings, compared to kernel-based distribution metrics such as Maximum Mean Discrepancy (MMD) (Wang et al., 2023).
Problem setup. We aim to minimize layer-wise MMC shifts under cross-modality FSL settings. This poses several challenges that motivate the design of SRR pipeline and LWS control: (i) cross-modality transfer must handle structural differences (e.g., input formats and feature shapes) and semantic discrepancies (e.g., non-overlapping label spaces), requiring an objective defined purely on latent representations (e.g., activation-based MMC statistics) rather than shared semantic spaces or paired data; (ii) the latent feature distribution misalignment is complicated by noisy representations and strong layer-wise dependencies inherited from pre-training, making few-shot repair brittle; and (iii) not all layers remain useful after repairing (Figure 2(c)), selectively recombining useful layers is critical and must be done efficiently and stably in multi-source settings with large candidate pools. Figure 3 overviews XTransfer.
Figure 3: Overview. XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control). LWS control first segments source models into layers and operates layer-wise search across pools. At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest. These layers are incrementally recombined during the search, restructuring models for enabling human sensing at the edge. Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.